OpenStudy (anonymous):

how do i find the value of k with the piecewise function k=x^2-2 when xis less than or equal to 0 and 3x+k when x is more than 0

OpenStudy (phi):

is there a typo in the question? Is it really k = x^2-2 (and not y=x^2-2 ?)